OpenStudy (anonymous):

Which method would be the simplest to solve. 7x+5y=19 -7x-2y=-16 A. graphing B. substitution C. elimination D. distribution

OpenStudy (anonymous):

2. what is the solution? 7x+5y=19 -7x-2y=-16 A.(2,1) B.(1,2) C.(7,5) D.(-7,-2)

OpenStudy (anonymous):

Ill give a medal

OpenStudy (btaylor):

1. Since you have +7x in one equation and -7x in the other, I'd use elimination. You end up with 3y=3, so y=1. A is your answer.

OpenStudy (anonymous):

thank you